The physical memory is split into physical pages (aka page frames), which have equal size. We know that the number of page frames in the physical memory equals the number of bytes in a page frame. We also know that physical addresses have 12 bits. How many bytes are in a page frame?
Answer #1

Physical address size is 12 bits;

p = number of page frames

b = number of bytes in a page frame

We are given that physical addresses have 12 bits

A = p + b = 12 bits

We are given that the number of page frames in the physical memory equals the number of bytes in a page frame.that means

p = b

Hence,

A = p + b

12 = p + p

12 = 2p

p = 12/2

p = 6

We know that p=b hence...

p = b = 6 bits

page frame size \large =2^{b}=2^{6}=64 bytes